A husband and wife work different shifts at the same plant. When the husband leaves from work to make the 20 mile trip home, the wife leaves their home and drives to work. They travel on the same road. The husband's driving rate is 45mph and the wife's is 35mph. How long into their drives can the wave at each other when passing on the road? Please make the chart as well.
Answer by stanbon(75887)

A husband and wife work different shifts at the same plant. When the husband leaves from work to make the 20 mile trip home, the wife leaves their home and drives to work. They travel on the same road. The husband's driving rate is 45mph and the wife's is 35mph. How long into their drives can the wave at each other when passing on the road? Please make the chart as well.
----
Husband DATA:
distance = x miles: rate = 45 mph ; time = d/r = x/45 hrs
-----
Wife DATA:
distance = 20-x miles ; rate = 35 mph ; time = d/r = (20-x)/35 hrs
------
Equation:
time = time
x/45 = (20-x)/35
----
35x = 45*20 - 45x
80x = 45*20
x = (1/4)(45) = 11 1/4 miles
-----
time = x/45 = (45/4)/45 = 1/4 hour
------------------
Cheers,
Stan H.
-------------------